# Some Optimizers are More Equal: Understanding the Role of Optimizers in Group Fairness

### v1

@@ Appendix E Proof of Lemma 1
where $z$ denotes the subgroup, and $\hat{y}$ represents the output of a model. The gap in demographic parity serves as a fairness measure [10.1145/3457607]. A common approach to quantifying this gap is through the zero-one loss difference [li2022achieving]: $|[\ell_{0/1}(w)|z=0]-[\ell_{0/1}(w)|z=1]|$. As established in prior work [li2022achieving], the zero-one loss can be approximated by the training loss, leading to the following fairness measure $\mathcal{F}(w)=|[\mathcal{L}(w)|z=0]-[\mathcal{L}(w)|z=1]|$. Substituting the given loss functions $\mathcal{L}_{0}(w)$ and $\mathcal{L}_{1}(w)$, we obtain $\mathcal{F}(w)=2|w|$. Since $\mathcal{F}(w)$ achieves its minimum at $w=0$, this proves that $w=0$ is the fairest minimizer.
Appendix F Training protocol
For facial expression recognition, we employ RetinaFace [deng2020retinaface] to detect and crop faces from images. The images in the FairFace dataset are already cropped by default. As part of the preprocessing pipeline for all tasks, we normalize the images and resize them to $224\times 224$. Data augmentation consists of random horizontal flipping and random rotation. Facial expression recognition and gender classification tasks are trained using the cross-entropy loss $\mathcal{L}_{ce}$, defined as:
For the multi-label classification task, given the class imbalance in MS-COCO, we employ the focal loss [lin2017focal] to mitigate overfitting to majority classes. It is formulated as:
where $\xi$ is the balancing factor and $\upsilon$ is the focusing parameter. To optimize the learning rate during training, we leverage Bayesian optimization [NIPS2012_05311655] combined with the Hyperband algorithm [li2018hyperband]. The best-performing models are trained twice, and we report the average performance across these runs.
We experiment with multiple backbone architectures, including ResNet-18 and ResNet-50 [he2016deep], VGG-16 [simonyan2014very], and a recently proposed vision transformer (tiny ViT)[wu2022tinyvit]. These architectures have been widely used in studies analyzing the impact of optimization methods on different learning paradigms[maunderstanding], as well as in fairness-focused works [10.1007/978-3-030-65414-6_35]. ResNet models leverage residual connections to facilitate deep representation learning, while VGG-16 provides a structured CNN design with uniform filter sizes. The tiny ViT, a transformer-based model, processes images as patch embeddings, and showed state-of-the-art performance [wu2022tinyvit]. This diverse selection allows us to investigate fairness effects across both convolutional and transformer-based architectures.
Appendix G Fairness and performance measures
We report the performance across different tasks using accuracy and F1. We evaluate group fairness using three widely recognized fairness criteria [10.1145/3457607]: equalized odds, equal opportunity, and demographic parity. A predictor is fair under equalized odds definition if the true positive and false positive rates are equal across demographic subgroups. We measure the violation of this criterion as:
$F_{EOP}=\min_{i,j}(\frac{\sum_{c=1}^{C}\text{{TPR}}_{i,c}}{\sum_{c=1}^{C}\text{{TPR}}_{j,c}}).$ (47)
Where $\text{{TPR}}_{i,c}$ denotes the true positive rate for i-th subgroup on class $c$. Finally, a predictor is said to be fair from demographic parity point of view if the predicted label distribution is independent of sensitive attributes. The corresponding measure is defined as:
$F_{DPA}=\min_{i,j,q}\frac{p(\hat{y}=c_{i}|z_{j})}{p(\hat{y}=c_{i}|z_{q})}.$ (48)
Regarding all three fairness metrics we use, higher values indicate better (fairer) outcomes.
Appendix H Additional results for Theorem 1
@@
Appendix I Datasets
@@
CelebA [liu2015faceattributes] is a large-scale and real-world dataset, which include 202,599 facial images from over 10,000 individuals. Each image is annotated with 40 distinct attributes, and the dataset is officially partitioned into training, validation, and test sets. In this study, we use gender (male vs. female) and age (young vs. old) as sensitive attributes, while performing facial expression recognition based on the “happy” attribute. Notably, the dataset exhibits an imbalance in demographic distribution, with the probability of an image depicting a male being 0.42 and the probability of depicting a young individual being 0.77. This skewed distribution indicates the presence of bias in the training set.
@@
FairFace [karkkainenfairface] is a large-scale face attribute dataset, mainly designed to mitigate racial bias in facial analysis tasks. It includes 108,501 facial images sourced primarily from the YFCC-100M Flickr dataset [thomee2016yfcc100m]. Each image is annotated for gender, age (9 sub-groups), and race (7 sub-groups). In this study, we leverage FairFace for gender classification, treating age and race as sensitive attributes. Given its broad demographic representation, FairFace serves as a robust benchmark in fairness related tasks.
@@
MS-COCO [10.1007/978-3-319-10602-1_48] is a large-scale, real-world benchmark dataset widely used for multi-label classification, object detection, and image captioning. It consists of over 330,000 images, including more than 200,000 labeled images spanning 80 object categories. In this study, we leverage MS-COCO for multi-label classification. Notably, the dataset does not provide explicit gender annotations. To infer gender labels, we follow a protocol established in previous works [zhao2017men], utilizing the five associated captions per image. Specifically, we retain images where at least one caption contains either the term “man” or “woman”. Additionally, we remove object categories that are not strongly associated with humans. Human-associated objects are defined as those appearing at least 100 times alongside either “man” or “woman” in the captions, resulting in a reduced set of 75 object categories. Fig. 10 provides sample captions and their corresponding inferred gender labels. It is worth mentioning that the probability of an image depicting a male is 0.71.
@@
Appendix J Hyperparameters
We employ Bayesian Optimization [NIPS2012_05311655] with Hyperband [li2018hyperband] algorithm to systematically tune the learning rate across different optimizers. Hyperband efficiently allocates training resources to more promising setups and progressively eliminate underperforming ones. It evaluates training at specific checkpoints, called “brackets”. We set Hyperband brackets to [10, 20, 40], allowing enough time for each run to demonstrate its potential. Additionally, the learning rate is sampled using a log-uniform distribution within the following search spaces:
  • SGD: $[10^{-3},10^{-1}]$
  • RMSProp: $[10^{-5},10^{-2}]$
  • Adam: $[10^{-5},10^{-2}]$
@@
Appendix K Additional fairness results across different backbones and datasets
@@
Appendix L Additional classification results across different backbones and datasets

### v2

@@ Appendix E Proof of Lemma 1
where $z$ denotes the subgroup, and $\hat{y}$ represents the output of a model. The gap in demographic parity serves as a fairness measure [38]. A common approach to quantifying this gap is through the zero-one loss difference [24]: $|[\ell_{0/1}(w)|z=0]-[\ell_{0/1}(w)|z=1]|$. As established in prior work [24], the zero-one loss can be approximated by the training loss, leading to the following fairness measure $\mathcal{F}(w)=|[\mathcal{L}(w)|z=0]-[\mathcal{L}(w)|z=1]|$. Substituting the given loss functions $\mathcal{L}_{0}(w)$ and $\mathcal{L}_{1}(w)$, we obtain $\mathcal{F}(w)=2|w|$. Since $\mathcal{F}(w)$ achieves its minimum at $w=0$, this proves that $w=0$ is the fairest minimizer.
Appendix F Additional Results for Theorem 1
@@
Appendix G Extension of Theorem 2 to Anisotropic Noise
Theorem 4.
Consider a population that consists of two subgroups with subgroup-specific loss functions $\mathcal{L}_{0}(w)$ and $\mathcal{L}_{1}(w)$, sampled with probabilities $p_{0}$ and $p_{1}$, respectively. Suppose a stochastic (online) training regime, in which each parameter update is computed from a sample drawn from one of the two subgroups. Suppose the gradients $\nabla\mathcal{L}_{0}(w_{k})$ and $\nabla\mathcal{L}_{1}(w_{k})$ are well-behaved anisotropic NGOs, namely $\mathcal{N}(\mu_{0},\Sigma_{0})$ and $\mathcal{N}(\mu_{1},\Sigma_{1})$, respectively. Then, the difference in parameter updates between subgroups 0 and 1 under RMSProp has an upper bound given by the corresponding difference under SGD. Consequently, RMSProp offers fairer updates across subgroups.
Proof.
The proof is similar to Theorem 2. For SGD, the difference between the parameter updates across subgroups is $\nabla\mathcal{L}_{0}-\nabla\mathcal{L}_{1}||$. Regarding RMSProp, let’s compute the expected value and the variance of $v_{k}$:
$\mathbb{E}[v_{k}]=\gamma^{k}v_{0}+(1-\gamma^{k})[p_{0}\left(\mu_{0}^{2}+diag(\Sigma_{0})\right)+p_{1}\left(\mu_{1}^{2}+diag(\Sigma_{1})\right)].$ (42)
Where $diag(\Sigma_{0})$ and $diag(\Sigma_{1})$ are the vectors with the diagonal elements of $\Sigma_{0}$ and $\Sigma_{1}$, respectively. Similar to the the isotropic case, we can show that the variance of $v_{k}$ is negligible and thus $v_{k}\rightarrow E[v_{k}]$. Hence, the difference between parameter updates are $D(\nabla\mathcal{L}_{0}-\nabla\mathcal{L}_{1})||$, where $D$ is a diagonal matrix, whose $jj$th element can be computed as:
$D_{jj}=\frac{1}{\sqrt{p_{0}\left(\mu_{0_{j}}^{2}+\sigma_{0_{j}}^{2}\right)+p_{1}\left(\mu_{1_{j}}^{2}+\sigma_{1_{j}}^{2}\right)+\epsilon}}.$ (43)
Prior work [37] has experimentally shown that $\sigma^{2}>\mu^{2}$. Additionally, in stochastic (online) training setup: $\Theta^{2}\rightarrow 1$. Thus, $D_{jj}<1$. This completes the proof. ∎
Appendix H Proof of Theorem 3
Proof.
We can compute $\nabla\mathcal{F}(w_{k})$ as:
$\nabla\mathcal{F}(w_{k})=\varphi(w_{k})\cdot(\nabla\Psi(w_{k})).$ (45)
Similarity, for RMSProp, we have:
In the equation above, as diagonal elements of $D$ are strictly less than 1, we have a strict inequality. Comparing Eqs. 48 and 49, we conclude the proof. ∎
Appendix I Extension of Theorem 3 to Anisotropic Noise
Theorem 5.
Consider a population that consists of two subgroups with subgroup-specific loss functions $\mathcal{L}_{0}(w)$ and $\mathcal{L}_{1}(w)$, sampled with probabilities $p_{0}$ and $p_{1}$, respectively. Suppose a stochastic (online) training regime in which each parameter update is computed from a sample drawn from one of the two subgroups. Suppose that the gradients $\nabla\mathcal{L}_{0}(w_{k})$ and $\nabla\mathcal{L}_{1}(w_{k})$ are well-behaved anisotropic NGOs, namely $\mathcal{N}(\mu_{0},\Sigma_{0})$ and $\mathcal{N}(\mu_{1},\Sigma_{1})$, respectively, with $\mu_{1}$ > $\mu_{0}$. Then, in expectation, the worst-case increase in the demographic parity gap after one iteration of RMSProp has an upper-bound no greater than the corresponding increase under SGD.
Proof.
The proof follows directly as the proof of this Theorem relies on the conclusion of the extension of the Theorem 2, which is already proved in Theorem 4. ∎
Appendix J Datasets
@@ Appendix I Extension of Theorem 3 to Anisotropic Noise
CelebA [31] is a large-scale and real-world dataset, which include 202,599 facial images from over 10,000 individuals. Each image is annotated with 40 distinct attributes, and the dataset is officially partitioned into training, validation, and test sets. In this study, we use gender (male vs. female) and age (young vs. old) as sensitive attributes, while performing facial expression recognition based on the “happy” attribute. Notably, the dataset exhibits an imbalance in demographic distribution, with the probability of an image depicting a male being 0.42 and the probability of depicting a young individual being 0.77. This skewed distribution indicates the presence of bias in the training set.
@@
FairFace [16] is a large-scale face attribute dataset, mainly designed to mitigate racial bias in facial analysis tasks. It includes 108,501 facial images sourced primarily from the YFCC-100M Flickr dataset [53]. Each image is annotated for gender, age (9 sub-groups), and race (7 sub-groups). In this study, we leverage FairFace for gender classification, treating age and race as sensitive attributes. Given its broad demographic representation, FairFace serves as a robust benchmark in fairness related tasks.
@@
MS-COCO [27] is a large-scale, real-world benchmark dataset widely used for multi-label classification, object detection, and image captioning. It consists of over 330,000 images, including more than 200,000 labeled images spanning 80 object categories. In this study, we leverage MS-COCO for multi-label classification. Notably, the dataset does not provide explicit gender annotations. To infer gender labels, we follow a protocol established in previous works [57], utilizing the five associated captions per image. Specifically, we retain images where at least one caption contains either the term “man” or “woman”. Additionally, we remove object categories that are not strongly associated with humans. Human-associated objects are defined as those appearing at least 100 times alongside either “man” or “woman” in the captions, resulting in a reduced set of 75 object categories. Fig. 10 provides sample captions and their corresponding inferred gender labels. It is worth mentioning that the probability of an image depicting a male is 0.71.
@@
Appendix K Training Protocol
For facial expression recognition, we employ RetinaFace [8] to detect and crop faces from images. The images in the FairFace dataset are already cropped by default. As part of the preprocessing pipeline for all tasks, we normalize the images and resize them to $224\times 224$. Data augmentation consists of random horizontal flipping and random rotation. Facial expression recognition and gender classification tasks are trained using the cross-entropy loss $\mathcal{L}_{ce}$, defined as:
For the multi-label classification task, given the class imbalance in MS-COCO, we employ the focal loss [28] to mitigate overfitting to majority classes. It is formulated as:
where $\xi$ is the balancing factor and $\upsilon$ is the focusing parameter. To optimize the learning rate during training, we leverage Bayesian optimization [50] combined with the Hyperband algorithm [23]. The best-performing models are trained twice, and we report the average performance across these runs.
We experiment with multiple backbone architectures, including ResNet-18 and ResNet-50 [11], VGG-16 [49], and a recently proposed vision transformer (tiny ViT)[54]. These architectures have been widely used in studies analyzing the impact of optimization methods on different learning paradigms[34], as well as in fairness-focused works [55]. ResNet models leverage residual connections to facilitate deep representation learning, while VGG-16 provides a structured CNN design with uniform filter sizes. The tiny ViT, a transformer-based model, processes images as patch embeddings, and showed state-of-the-art performance [54]. This diverse selection allows us to investigate fairness effects across both convolutional and transformer-based architectures.
Appendix L Hyperparameters
We employ Bayesian Optimization [50] with Hyperband [23] algorithm to systematically tune the learning rate across different optimizers. Hyperband efficiently allocates training resources to more promising setups and progressively eliminate underperforming ones. It evaluates training at specific checkpoints, called “brackets”. We set Hyperband brackets to [10, 20, 40], allowing enough time for each run to demonstrate its potential. Additionally, the learning rate is sampled using a log-uniform distribution within the following search spaces:
  • SGD: $10^{-3},10^{-1}]$
  • RMSProp: $10^{-5},10^{-2}]$
  • Adam: $10^{-5},10^{-2}]$
@@
Appendix M Additional Fairness Results across Different Backbones and Datasets
@@
Appendix N Additional Fairness Results across Different Variants of Optimizers
Our theoretical analysis identifies gradient normalization via second moment as the key mechanism for improving group fairness. This provides a clear framework for predicting the behavior of other optimizers with regards to fairness. To validate these predictions, we have conducted a new set of experiments with SGD w/ momentum [30], AdamW [32], and AdaBound [33] variants. Following the protocol from our main experiments, we use the CelebA dataset with a ViT backbone, treating gender and age as sensitive attributes. We tuned the learning rate for each optimizer using Bayesian optimization with Hyperband and report the average fairness scores in Table 6. As shown in the table, AdamW consistently achieves the best fairness scores. This is expected, as AdamW retains the normalization mechanism of Adam and RMSProp via the second moment. In contrast, SGD w/momentum, lacking the normalization via the second moment, performs more biased. The performance of AdaBound is comparable to SGD w/momentum. The reason is that the AdaBound algorithm deliberately clips the second moments to make it act like SGD, which in turn limits the fairness advantage.
Table 6: Fairness comparison of optimizer variants.
Gender Age
AdamW AdaBound SGD w/ mom. AdamW AdaBound SGD w/ mom.
$F_{\text{EOD}}$ 64.90 61.10 61.12 71.11 67.20 67.25
$F_{\text{EOP}}$ 99.90 96.10 95.07 98.95 96.60 96.55
$F_{\text{DPA}}$ 73.11 70.01 69.50 98.11 95.39 95.39
Appendix O Additional Classification Results across Different Backbones and Datasets